# Robust Prompt Optimization for Defending Language Models Against Jailbreaking Attacks

### v3

@@ 4.3 Ablations
Adaptive attack results. We also consider a more challenging threat model where the adversary white-box access to the defense. For the perplexity filter defense, we use the perplexity-regularized variant of GCG Jain et al. (2023), which modifies the optimization objective to reduce perplexity. SmoothLLM and the rephrasing defense are not straightforward to attack in this setting, so we use the generic adversary. For RPO, we consider an adversary that can optimize an attack on top of the RPO suffix. We find in Tab. 4 that RPO retains state-of-the-art performance on JailbreakBench even under this more challenging threat model. We observe a small 4% ASR increase on PAIR and a 1% ASR increase on GCG for Vicuna. Notably, Llama-2 retains full robustness to both PAIR and GCG. Generally, we observe that optimizing a GCG string on an RPO suffix for 500 steps cannot break it, while the adaptive PAIR attack can induce affirmative responses that are not harmful or irrelevant.
Effect on benign use and cost. As LLMs become increasingly deployed in real-world contexts, it is imperative for defenses to be practical, cheap, and not greatly affect benign use. In Tab. 4, we evaluate models with RPO on MT-Bench Zheng et al. (2023), which evaluates multiturn interaction and Multitask Language Understanding (MMLU) Hendrycks et al. (2021), which evaluates domain knowledge. We observe that MMLU performance is largely unaffected, but we observe a small performance reduction on MT-Bench. Surprisingly, we find that using RPO on benign prompts will not cause unnecessary rejection even when optimizing the suffix on only harmful instructions. This may be due to the implicit tendency of models to reject these instructions, suggesting RPO only strengthens this inclination rather than introducing a new refusal mechanism. The responses also do not appear qualitatively different, except in some cases where the model explicitly mentions the suffix in its response. This failure case occurs mainly with shorter instructions. An example of this is provided in Sec. C in the Appendix. To minimize this effect for production, we suggest only applying RPO to longer queries, as stronger attacks also increase input length. This could also be mitigated by optimizing semantically meaningful suffixes, which we leave to future work.
Additionally, RPO has a small inference cost of 20 additional tokens. This is generally much lower than baselines such as SmoothLLM, which involve at least twice as many additional full inference calls as normal usage. Finally, while RPO suffixes can be expensive to optimize, this can be offset by the ease of transferability to other models compared to other defenses with high computational costs, such as adversarial training, which is local to the base model.
@@ 5 Limitations and Conclusion
In this paper, we propose Robust Prompt Optimization (RPO), an approach for improving the robustness of LLMs against jailbreaking attacks. By formalizing an optimization-based objective that directly incorporates the threat model, RPO generates transferable and lightweight defensive suffixes that are robust to a wide range of attacks, including unseen ones. Our experiments on JailbreakBench and HarmBench demonstrate RPO’s superior performance compared to existing defenses, reducing attack success rates significantly across different models while incurring only minor effects on benign usage. This suggests text-based jailbreaking may be an easier problem to address than adversarial attacks in vision. However, RPO does not currently cover multimodal models or other failure modes such as deception and malicious code generation. Proposing our defense may also lead to the development of stronger attacks, including those that can break RPO. Indeed, while we observe high transferability to new attacks, using RPO does not typically result in full robustness. Future directions include optimizing defenses on a greater variety of attacks, combining various defenses into comprehensive guardrails, and red-teaming to discover new security risks.
@@ 6 Acknowledgements
We thank Mantas Mazeika and Yi Zeng for their helpful discussions. This research used Illinois Compute to provide computing and data resources.
@@ A.1 Experiment Details
We follow the attack and defense settings provided in JailbreakBench and Harmbench for our baselines. We use the test cases and jailbreak prompts provided in each benchmark. For GCG, the adversarial suffix is optimized for each individual instruction, for a batch size of 512 and 500 optimization steps. This setting is followed in our adaptive attack setup. For closed-source models, these suffixes are optimized on an open-source model and directly transferred. For PAIR, the attacker model is Mixtral Jiang et al. [2024] with a temperature of one, top-p sampling with p = 0.9, N = 30 streams, and a maximum depth of K = 3. This setting is also used in our adaptive attack setup. For JBC, we use the most popular jailbreak template, named “Always Intelligent and Machiavellian” (AIM).

### v5

@@ 4.3 Ablations
Adaptive attack results. We also consider a more challenging threat model where the adversary white-box access to the defense. For the perplexity filter defense, we use the perplexity-regularized variant of GCG (Jain et al., 2023), which modifies the optimization objective to reduce perplexity. SmoothLLM and the rephrasing defense are not straightforward to attack in this setting, so we use the generic adversary. For RPO, we consider an adversary that can optimize an attack on top of the RPO suffix. We find in Tab. 4 that RPO retains state-of-the-art performance on JailbreakBench even under this more challenging threat model. We observe a small 4% ASR increase on PAIR and a 1% ASR increase on GCG for Vicuna. Notably, Llama-2 retains full robustness to both PAIR and GCG. Generally, we observe that optimizing a GCG string on an RPO suffix for 500 steps cannot break it, while the adaptive PAIR attack can induce affirmative responses that are not harmful or irrelevant.
Effect on benign use and cost. As LLMs become increasingly deployed in real-world contexts, it is imperative for defenses to be practical, cheap, and not greatly affect benign use. In Tab. 4, we evaluate models with RPO on MT-Bench (Zheng et al., 2023), which evaluates multiturn interaction and Multitask Language Understanding (MMLU) (Hendrycks et al., 2021), which evaluates domain knowledge. We observe that MMLU performance is largely unaffected, but we observe a small performance reduction on MT-Bench. Surprisingly, we find that using RPO on benign prompts will not cause unnecessary rejection even when optimizing the suffix on only harmful instructions. This may be due to the implicit tendency of models to reject these instructions, suggesting RPO only strengthens this inclination rather than introducing a new refusal mechanism. The responses also do not appear qualitatively different, except in some cases where the model explicitly mentions the suffix in its response. This failure case occurs mainly with shorter instructions. An example of this is provided in Sec. C in the Appendix. To minimize this effect for production, we suggest only applying RPO to longer queries, as stronger attacks also increase input length. This could also be mitigated by optimizing semantically meaningful suffixes, which we leave to future work.
Additionally, RPO has a small inference cost of 20 additional tokens. This is generally much lower than baselines such as SmoothLLM, which involve at least twice as many additional full inference calls as normal usage. Finally, while RPO suffixes can be expensive to optimize, this can be offset by the ease of transferability to other models compared to other defenses with high computational costs, such as adversarial training, which is local to the base model. In practice, we find that optimizing an RPO suffix is around 8x cheaper than a GCG suffix and only takes a few hours, due to the natural tendency of LLMs to already refuse harmful instructions. The lower optimization cost also suggests RPO suffixes improve robustness by enhancing refusal mechanisms, allowing transfer to different LLMs and attacks.
@@ 5 Limitations and Conclusion
In this paper, we propose Robust Prompt Optimization (RPO), an approach for improving the robustness of LLMs against jailbreaking attacks. By formalizing an optimization-based objective that directly incorporates the threat model, RPO generates transferable and lightweight defensive suffixes that are robust to a wide range of attacks, including unseen ones. Our experiments on JailbreakBench and HarmBench demonstrate RPO’s superior performance compared to existing defenses, reducing attack success rates significantly across different models while incurring only minor effects on benign usage. This suggests text-based jailbreaking may be an easier problem to address than adversarial attacks in vision. However, RPO does not currently cover multimodal models, LLM agents, or other failure modes such as deception and malicious code generation. Proposing our defense may also lead to the development of stronger attacks, including those that can break RPO. Indeed, while we observe high transferability to new attacks, using RPO does not typically result in full robustness. Future directions include optimizing defenses on a greater variety of attacks, combining various defenses into comprehensive guardrails, and stronger red teaming strategies to discover new security risks.
@@ 6 Acknowledgements
We thank Mantas Mazeika and Yi Zeng for their helpful discussions and assistance with HarmBench. This work used NVIDIA GPUs at NCSA Delta through allocations CIS230218 and CIS230365 from the ACCESS program and from the Illinois Compute Campus Cluster.
@@ A.1 Experiment Details
We follow the attack and defense settings provided in JailbreakBench and Harmbench for our baselines. We use the test cases and jailbreak prompts provided in each benchmark. For GCG, the adversarial suffix is optimized for each individual instruction, for a batch size of 512 and 500 optimization steps. This setting is followed in our adaptive attack setup. For closed-source models, these suffixes are optimized on an open-source model and directly transferred. For PAIR, the attacker model is Mixtral [Jiang et al., 2024] with a temperature of one, top-p sampling with p = 0.9, N = 30 streams, and a maximum depth of K = 3. This setting is also used in our adaptive attack setup. For JBC, we use the most popular jailbreak template, named “Always Intelligent and Machiavellian” (AIM).